OpenStudy (anonymous):

Use the quadratic formula to solve the equation below. -6x= -4x^2+18 A.X= 3 | x= -15 B.X= 1.5 | C=-3 C.X= 3|X=-3 D.X= -3 | X= -1.5

OpenStudy (anonymous):

\[x= \frac{- ( 6 )^2 \pm \sqrt{ (-6)^2 - ( 4 )( 4 ) ( -18)} }{ 2 ( 4 )} \]

OpenStudy (anonymous):

Can you please explain this to me? @Luis_Rivera

OpenStudy (anonymous):

\[x= \frac{ 36\pm \sqrt{ -92 } }{8} \]

OpenStudy (anonymous):

now find x

OpenStudy (anonymous):

I don't understand how to do this. I got 2.6 the way I did it. Please explain

OpenStudy (anonymous):

@Luis_Rivera

hartnn (hartnn):

first, try to bring this -6x= -4x^2+18 in the form ax^2+bx+c=0 can you ?

OpenStudy (anonymous):

So I would move the equal sign to the end and leave to -6x^2 where it originally was but add the equal sign?

hartnn (hartnn):

i am not sure i understand that fully, but tell me what equation you got ?? i'll verify it.

OpenStudy (anonymous):

-6x^2+-4x^2+18=0

OpenStudy (anonymous):

Its probably wrong

hartnn (hartnn):

firstly, its 6x and not 6x^2 , right ? also, if you want to move -6x to right, you add 6x on both sides, so that on left you get -6x+6x =0 got this ?

OpenStudy (anonymous):

Oh yeah sorry -6x

OpenStudy (anonymous):

Yeah

hartnn (hartnn):

now will you please try again and tell me what equation you get ?

OpenStudy (anonymous):

-4x^2+18=0 Because we took out the -6x But would I add 6x to the 18?

hartnn (hartnn):

you would add 6x on right side too, right ? so -4x^2+6x+18 = 0

OpenStudy (anonymous):

Okay. So then would we us pemdas or a different step?

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

hartnn (hartnn):

tell me what you get for a,b,c, first

OpenStudy (anonymous):

I don't know how to find them. I never learned this.

OpenStudy (anonymous):

How would I find them?

hartnn (hartnn):

 now ?